OpenStudy (clamin):

Solve this equation by factoring::: 7x^2-21x-189= 7

OpenStudy (anonymous):

7(x^2-3x-27)=7 x^2-3x-27=1 x^2-3x-28 (x+4) (x-7) x= -4 , x= 7